OpenStudy (anonymous):

Solve: 2(x + 1) = −8

mathslover (mathslover):

\[\huge{2x+1=-8}\] \[\huge{2x=-9}\] \[\huge{x=-9/2}\]

mathslover (mathslover):

\( \begin{array}{} \cal{\large{\text{Hey! Good to see you on this site!}}}\\ \\ {\huge \mathtt{\color{blue}{\text{Welcome To OpenStudy!}}}} \\ \text{enjoy answering and asking questions here}\\ \cal{\large{\text{Tips:)}}} \\ \textbf{Ambassadors (people with A next to their name) are here to help you.}\\ \textbf{If you have any questions about the site, feel free to ask us!}\\ \textbf{Users who appear in Purple are Moderators and help run the site.}\\ \textbf{If you have a serious issue with the site or with the behavior of other}\\ \textbf{users}\\ \textbf{please feel free to raise it with them.}\\ \textbf{Want to study a particular subject? Click 'Find More Subjects'}\\ \textbf{search for it there.}\\ \textbf{If you get a good response from a user , click on best response}\\ \end{array}\) http://www.openstudy.com/code-of-conduct